OpenStudy (anonymous):

Can someone please tell me how to get the derivative of the square root of (1-x^2)

OpenStudy (anonymous):

you have to use what is called the chain rule: http://www.themathpage.com/acalc/chain.htm

OpenStudy (anonymous):

yes, i know that much

OpenStudy (anonymous):

see the link on how to solve your problem and post back if you have doubts.

OpenStudy (anonymous):

i looked at it but i don't get how to apply the chain rule to the problem... that is my problem

OpenStudy (anonymous):

You need to use the chain rule, which states that \[f(g(x)) = g \prime(x)*f \prime(g(x))\]. Setting f(x) = x^(1/2), and g(x) = 1-x^2, The formula yields \[-2x/\sqrt{1-x^2}\]. You may need to rationalize.

OpenStudy (anonymous):

srossd is right. that is the answer you should get.